# Comparing Quantum Annealing and Spiking Neuromorphic Computing for Sampling Binary Sparse Coding QUBO Problems

### v1

@@ 2.2 Implementation on Loihi
2.2.1 Lateral competition
Previous work has shown that sparse coding optimization problems, such as the one we focus on in eq. (1), can be solved using the dynamics of neural networks incorporating lateral inhibition, a biologically plausible implementation of a sparse solver referred to as a locally competitive algorithm (LCA). LCA aims to find a sparse solution of eq. (1) by allowing the dynamical system to evolve to a minimum energy configuration. However, such dynamical systems are susceptible to local minima.
2.2.2 Loihi 2 Neuromorphic Chip Implementation
The Intel Loihi 2 neuromorphic processor aims to mimic the neural structure and processing methods of the human, which can be used to address specific types of computational tasks. Compared to the previous generation of the Loihi processor, Loihi 2 has several expanded capabilities, namely improved manufacturing. Loihi 2 contains 1 million neurons per chip, and has about 120 million synapses.
Loihi 2 utilizes graded 8-bit spike events, which are different from the binary spike events used in Loihi 1. This change allows the processor to model neural information in a way that is closer to the biological functioning of the human brain. In particular, Loihi 1 enabled communication between artificial neurons with binary spikes. Those digital pulses represented the firing of a neuron by effectively transmitting a signal that has only two states: “on” or “off” (1 or 0). This binary approach mimics the all-or-nothing nature of neuron activity in biological brains at a fundamental level but with less granularity. With the introduction of graded 8-bit spike events instead of binary spikes, each spike can carry more nuanced information, with the intensity of the spike encoded in 8 bits.
In contrast with the Stochastic Constraint Integrate and Fire (SCIF) neuron model that was implemented in [henke2023sampling] for Loihi 1, using the new functionality of Loihi 2, we are able to deploy a Non-equilibrium Boltzmann (NEBM) neuron model specifically designed to address optimization problems such as the QUBO problem. This model is inspired by the principles of statistical mechanics and thermodynamics, particularly the concept of a non-equilibrium steady state in systems that are driven by external forces.
The NEBM neuron model diverges from traditional artificial neuron models by incorporating the dynamics of non-equilibrium systems. In typical computational models like the Boltzmann machine, neurons are treated as binary units that reach an equilibrium state where the probability distribution of their states follows the Boltzmann distribution. In contrast, the NEBM model operates under non-equilibrium conditions, meaning the system does not settle into a thermal equilibrium but is instead continuously driven by external inputs or energy sources. Unlike equilibrium models where state transitions depend solely on minimizing the system’s energy, NEBM incorporates dynamics where transitions also depend on external driving forces. This allows the system to explore more of the solution space and potentially escape local minima more effectively. In NEBM, energy is not conserved but dissipated due to the system’s open and driven nature. This aspect is crucial for implementing effective annealing processes, where the system can cool down and stabilize towards a solution of the optimization problem. The NEBM model also includes time as a critical factor, with the system’s state evolving over time rather than settling into a static equilibrium. This temporal aspect allows for dynamic updating rules that can be tuned for faster convergence or better exploration of the solution space.
For the task of optimizing a QUBO problem using the NEBM neuron model, the first step is to map each variable in the QUBO to a neuron in the NEBM. The interactions defined in the QUBO matrix are translated into synaptic connections between these neurons. Similar to simulated annealing, the NEBM model can implement a form of dynamic annealing where the system gradually stabilizes by reducing the impact of external drives and allowing the system to settle into a low-energy configuration that corresponds to an optimal or near-optimal solution of the QUBO problem. The system is initialized in a non-equilibrium state and evolves according to specified dynamics, driven by both internal neuronal computations and external inputs. These dynamics are designed to reduce the overall energy defined by the QUBO, but with considerations for non-equilibrium behavior that allows exploration. The NEBM approach can be programmed onto neuromorphic hardware like Loihi 2, which can mimic these dynamic and non-equilibrium processes efficiently, using spiking neural networks.
The exact hyperparameters used to run Loihi 2 are as follows. The neuron model is set to nebm-sa-refract. The annealing schedule is linear. The maximum temperature is $15$, and the minimum temperature is $1$, and the number of steps per temperature is $20$, with temperature delta set to $1$. The parameter refract_scaling is set to a random integer between 4 and 8, and refract is set to a random integer between $5$ and $10$. The refract_counter is set to 0.
@@ 2.3 Quantum Annealing Implementation
Quantum annealing is a type of analog quantum computation that generally has the goal of computing low energy states in a physical system that correspond to low energy states of a classical spin system [Kadowaki_1998, Farhi2000, Santoro2006, morita2008mathematical, das2008colloquium]. A review of quantum annealing can be found in [Hauke_2020]. D-Wave manufactures quantum annealers using superconducting flux qubits [Johnson2011, Lanting_2014], several generations of which have been utilized in a large number of studies of different optimization problems and physical spin systems [King_2022, harris2018phase, boixo2013experimental, Venturelli_2015, Boixo_2014].
The D-Wave quantum annealer is designed to minimize QUBO problems, given by functions of the form of eq. (4). In eq. (4), the linear terms $h_{i}\in\mathbb{R}$ and the quadratic couplers $Q_{ij}\in\mathbb{R}$ are chosen by the user to define the problem under investigation. The variables $a_{1},\ldots,a_{n}\in\mathbb{B}$ are binary and unknown. The aim is to find a configuration of $a_{1},\ldots,a_{n}$ which minimizes eq. (4).
Quantum annealing is based on adiabatic quantum computation, where the system is initialized in the ground state of a Hamiltonian that is easy to prepare, in the case of D-Wave this is the transverse field Hamiltonian. The system is then slowly transitioned to a Hamiltonian of interest, whose ground state is unknown (and, likely hard to find). In ideal adiabatic settings, namely if the transition is sufficiently slow, the system will remain in the ground state, leading to the ground state of the final Hamiltonian. Therefore, this can be used as a method of computing ground states of spin systems (Ising models). The system is described as
$H_{\text{system}}(t)=A(t)H_{\text{init}}+B(t)H_{\text{problem}},$ (5)
where $A(t)$ and $B(t)$ define the anneal schedule, $H_{\text{init}}$ is the transverse field Hamiltonian, and $H_{\text{problem}}$ is the problem Ising model that we wish to find the optimal solution of. Ising models can easily be converted into QUBOs, and vice versa, and since many optimization problems can be formulated as QUBOs, quantum annealing can be used to sample solutions of a large number of optimization problems [Lucas2014, Yarkoni2021].
@@
$H_{\text{system}}=-\frac{A(s)}{2}\Big{(}\sum_{i}^{n}\sigma^{x}_{i}\Big{)}+\frac{B(s)}{2}\Big{(}H_{\text{problem}}\Big{)},$ (6)
where $s\in[0,1]$ is a variable called the anneal fraction which defines a weighting of both the transverse field and problem Hamiltonians (by $A(s)$ and $B(s)$) at each point in time during the anneal, $\sigma^{x}_{i}$ is the Pauli matrix acting on each qubit $i$, $H_{\text{problem}}$ is the problem Ising model that we wish to sample low energy states from. In our case, $H_{\text{problem}}$ is the QUBO of eq. (4) that has been converted to an Ising model such that all of the variable states are spins. The user can program a variety of additional features on the device, including different anneal schedules and annealing times.
@@ 2.3.1 Parallel Quantum Annealing Minor Embeddings
In order to make more effective use of the large hardware graph of D-Wave’s Advantage_system4.1 (compared to the QUBO models which are being embedded), we employ parallel quantum annealing, also tiling, in which multiple Ising models can be embedded on the chip such that they are disjoint and are solved in parallel during a single anneal [Pelofske2022parallel, Pelofske2022solving, PhysRevA.91.042314]. As shown in Figure 6, we use multiple all-to-all minor embeddings onto the D-Wave hardware graph, thus allowing us to solve the same problem multiple times simultaneously during a single anneal readout cycle. This allows us to obtain more samples per anneal, but also mitigates local biases and errors that may be present on some parts of the chip (as well as differences in the random minor-embeddings). Moreover, parallel quantum annealing partially mitigates the effects of using a single sub-optimal embedding due to multiple different random embeddings being used, and also different parts of the hardware graph being used (since the D-Wave hardware has non-uniform noise properties across the different components). Note that parallel quantum annealing is also referred to as tiling in the D-Wave documentation  [^2] https://dwave-systemdocs.readthedocs.io/en/samplers/reference/composites/tiling.html. Throughout all experiments, we consider all-to-all embeddings for sizes $64$ and $128$ variables. Those minor embeddings can be computed with the heuristic tool minorminer [minorminer], applied iteratively to the hardware graph in order to obtain disjoint minor embeddings, using the methods described in [Pelofske2022parallel] which attempts to maximize the number of disjoint minor embeddings that can be embedded on the hardware graph.
@@
2.3.2 Choice of D-Wave parameters
@@
For all experiments involving reverse annealing, the schedule used is $\{[0,1],[10,s],[90,s],[100,1]\}$, where each tuple defines a point in time (from the start to the end of the anneal process) and an anneal fraction $s$. The anneal fraction is the normalized time used to control how the quantum Hamiltonian is moved from the initial superposition of states to the problem QUBO (which is a classical Hamiltonian) during the anneal [annealing]. The anneal schedule is constructed by linear interpolation between those four points. The reverse anneal schedule we use is symmetric with a pause of $80$ microseconds. It has an increasing and decreasing ramp on either side of a pause of duration $10$ microseconds. We vary the anneal fraction $s$ at which the pause occurs.
@@ 2.3.3 Quantum Evolution Monte Carlo with Reverse Annealing
We will use the D-Wave annealer in connection with a Monte Carlo chain of reverse anneals, in which the best solution of an initial set of anneal-readout cycles is encoded as the initial state of the next anneal, and then the best solutions found from each iteration are used as the initialized classical spin state of the subsequent (reverse) anneals. This process works on the logical problem, meaning after unembedding of all chained qubits.
@@
To be precise, a sequence of reverse anneals are chained together in a Monte Carlo like process, where each subsequent round of reverse anneals is initialized with a classical state that is defined by the best solution found at the last set of reverse anneals. This chain of reverse anneals is initialized with the best solution found from a $100$ microsecond forward anneal with $1000$ samples and the standard linear anneal schedule (forward anneal denotes that it is a standard annealing process, initialized in a uniform superposition and not in a specific configuration of spins). Because of the use of parallel quantum annealing, there is a multiplier on the total number of samples obtained at each QEMC step; for the $64$ variable QUBO problems we actually obtain $7,0000$ samples per $1,000$ sample anneal readout cycle, and for the $128$ variable QUBO problems we obtain $2,000$ samples per $1,000$ sample anneal readout cycle. Each reverse annealing step in the chain uses the parameter reinitialize_state=True when executing on the D-Wave quantum annealers, which re-initializes the state of the reverse anneal after each anneal-readout cycle. This technique has been used in other contexts, and is referred to as both “iterative reverse annealing” [Yamashiro2019, Bando2022, Arai2021, pelofske2023initial] and “quantum evolution Monte Carlo” (QEMC) [King2021magnetization, King2018, King2021scaling, Kairys2020, LopezBezanilla2023]. The technique we use here differs slightly from some of these approaches in that we do not use a single anneal-readout cycle at each step of the Monte Carlo chain process, instead we use a batch of anneals at each step and initialize the next anneal with the lowest energy solution found in the current batch of anneals.
@@
Figure 2: Top: The best (lowest energy) reconstruction found using D-Wave with standard forward annealing (with no parallel QA and a default annealing time of $20$ microseconds), reverse annealing QEMC, neuromorphic computing on Loihi 2, and iterated warm starting neuromorphic computing with Loihi 2 (in order). Below: Experimentally computed figure reconstructions are the best mean energy and best mean sparsity across the $16$ QUBO models (for the best parameter combination found for each device or technique).
@@
Figure 3: Iterated warm starting on results on Loihi 2 for $100$ iterations for each of the $16$ QUBO problems (each QUBO contains $64$ variables). The minimum energy sample from the converged simulation is plotted at each iteration step.
@@ 2.5 Exact QUBO solutions using CPLEX
In order to determine how well all of the evaluated heuristic methods worked at sampling the optimal solution(s) of the generated QUBO models, we used the classical CPLEX [CPLEX] solver which can solve the generated QUBOs exactly. To this end, we first transform each QUBO into a Mixed Integer Quadratic Programming (MIQP) problem, and solve those problems exactly using the CPLEX solver [CPLEX]. The CPLEX solver does not provide information on whether the QUBOs have multiple optimal solutions, or only $1$, however it is expected that in general these QUBOs will not have degenerate ground states. We leave further investigations on the ground state degeneracy of these binary sparse coding QUBO problems to future research.
@@ 3 Results
Figure 3 presents the iterated warm started learning on Loihi 2 for all $16$ of the $64$ decision variable QUBO models. Interestingly, this process of warm starting the simulation using the state found during the previous computation seems to converge fairly fast to a local minimum for most of the QUBO problems. Notably, $4$ of the plots show high variability of the learning curve for those QUBOs, whereas for most of the other QUBOs there is very little variability of the learning curve. Figure 3 also plots the minimum energies computed using CPLEX. These results are the first that we are aware of for using iterated warm starting on a spiking neuromorphic processor.
@@
Figure 4 shows the QEMC, or iterated reverse annealing, progression when sampling all of the $16$ sparse coding QUBOs using $100$ iterations and varying anneal fractions at which the pause occurred. At each QEMC step, the minimum energy (e.g., the objective function evaluation) sample is plotted. Figure 4 shows that there are some anneal fractions ($s$) that produced the desired monotonic decrease of minimum computed energy at each QEMC step, whereas other anneal fractions did not consistently improve on solution quality. Specifically, the anneal fraction of $s=0.44$ clearly introduced too much of the transverse field Hamiltonian into each simulation, resulting in a loss of information on the best previous obtained solution. On the other hand, the choice $s=0.6$ did not utilize enough of the quantum fluctuations provided by the transverse field Hamiltonian to search for lower energy states, although the improvement was always monotonic. The anneal fractions of $s=0.5$ and $0.55$ showed the best tradeoff in our experiments, allowing for reasonably good exploration of the problem search space. Figure 4 shows that the RA QEMC method was able to find the optimal solution for some, but not all of of the QUBO problems. This is examined in more detail in Table 1. Figure 4 plots the minimum energy computed by simulated annealing as black horizontal dashed lines, and Table 1 shows that those solutions are also optimal as verified by CPLEX. When executing the QEMC simulations of Figure 4, the initial state was chosen to be the best sample found from an initial standard linear schedule forward anneal; that initialization is then repeated with independent anneals for each run of QEMC, meaning that the initial states all have slightly different objective values. If there were multiple solutions with the same objective value, one was chosen arbitrarily as the initial vector.
@@
The ordering of the sub-plots (in terms of the QUBO problem being solved) in Figure 3 and Figure 4 is consistent between the two figures. Notably, all of the iterated reverse annealing plots of Figure 4 show very clear and consistent convergence behavior (e.g. for $s=0.5$), but the iterated warm starting on Loihi 2 does not have as clear convergence behavior as a function of the iteration index.
@@
Table 1 also reports the sparsity of the optimal bitstring computed by CPLEX. Simulated annealing always correctly found the optimal solution. Loihi 1 found the optimal solution in 00 out of the $16$ QUBOs, so the optimal solution count is not shown. The Loihi 1 results are shown for comparison against Loihi 2, and are taken from the simulations in [henke2023sampling], where the minimum energies are taken from the distribution of $2000$ samples computed across a varied set of device parameters. The minimum energies reported both from iterated reverse quantum annealing and iterated warm starting on Loihi 2 are the minimum energies found from the entirety of the simulation (which, for converged simulations is the last point found after the $100$ iterations). The reported minimum energies from Loihi 2 are taken from the distribution of $100$ samples obtained from the iterated warm starting procedure.
@@
Table 1 verifies that the trained sparse coding QUBOs indeed have solution vectors that are reasonably sparse. However, although unlikely to occur for our types of QUBO models, there could exist multiple variable assignments that give the same optimal objective function value (i.e., degenerate ground states). Moreover, as CPLEX only returns one optimal solution, the reported CPLEX results do not capture the case of more than one optimal solution with different sparsity.
@@
Figure 4: Quantum Evolution Monte Carlo with reverse annealing, where the results for each of the $16$ QUBO problems (each QUBO contains $64$ variables) is shown by each sub-figure. The anneal fraction at which the reverse anneal is paused is varied, and the resulting minimum energy found in the $1000$ samples at each Monte Carlo step is plotted. All quantum annealing computations used an annealing time of $100$ microseconds.

### v2

@@ 2.3.6 Implementation on Loihi
The exact hyperparameters available for use to run Loihi 2 are as follows. The neuron model is set to nebm-sa-refract. The annealing schedule is linear. The maximum temperature ($T_{\max}$) is $15$, and the number of steps per temperature is $20$, with temperature delta ($\Delta_{T}$) set to $1$. The parameter refract_scaling is set to a random integer between 4 and 8, and refract is set to a random integer between $5$ and $10$. The refract_counter is set to 0. Each simulation is run for a total of 6000 time steps and sampled every 20 time steps, for a total of 300 samples per run.
@@
Despite Loihi 2’s advancements, several factors can lead to non-ideal performance when solving QUBO problems with the NEBM model. First, hardware constraints such as limited on-chip memory and communication bandwidth can restrict the size or connectivity of the problem that can be efficiently mapped, which may result in suboptimal partitioning or approximation of the underlying QUBO. Second, the discrete nature of spiking dynamics and the complexity of tuning neuron and synapse parameters mean that finding the right hyperparameters (e.g., annealing schedules, refractory settings) often involves extensive trial and error, leading to potential inefficiencies. Third, real-time spike communication overhead—particularly if signals must be routed across multiple cores or chips—can slow the effective convergence rate. Fourth, noise and variability inherent in spiking hardware may occasionally cause the system to settle in unfavorable local minima, especially for large-scale or high-dimensional problems. Finally, while non-equilibrium approaches can explore more of the solution space, they also rely on carefully calibrated driving forces, and any inaccuracies or mismatches in these driving inputs can impede overall performance. These factors, taken together, highlight that while Loihi 2 shows promise for neuromorphic optimization, further refinements in hardware design, software tooling, and parameter-tuning methodologies are critical to unlocking its full potential.
@@ 2.4 Quantum Annealing Implementation
Quantum annealing is a type of analog quantum computation that generally has the goal of computing low energy states in a physical system that correspond to low energy states of a classical spin system (i.e., diagonal Hamiltonian) [Kadowaki_1998, Farhi2000, Santoro2006, morita2008mathematical, das2008colloquium]. A review of quantum annealing can be found in [Hauke_2020]. D-Wave manufactures quantum annealers using superconducting flux qubits [Johnson2011, Lanting_2014], several generations of which have been utilized in a large number of studies of different optimization problems and physical spin systems [King_2022, harris2018phase, boixo2013experimental, Venturelli_2015, Boixo_2014].
The D-Wave quantum annealer is designed to minimize QUBO problems, given by functions of the form of eq. (5). In eq. (5), the linear terms $h_{i}\in\mathbb{R}$ and the quadratic couplers $Q_{ij}\in\mathbb{R}$ are chosen by the user to define the problem under investigation. The variables $a_{1},\ldots,a_{p}\in\mathbb{B}$ are binary and unknown. The aim is to find a configuration of $a_{1},\ldots,a_{p}$ which minimizes eq. (5).
Quantum annealing is based on adiabatic quantum computation [Kadowaki_1998, Farhi2000, Santoro2006, morita2008mathematical, das2008colloquium], where the system is initialized in the ground state of a Hamiltonian that is easy to prepare, in the case of D-Wave this is the Transverse field Hamiltonian. The system is then slowly transitioned to a (diagonal) Hamiltonian of interest, whose ground state is unknown (and, likely hard to find). In ideal adiabatic settings, namely if the transition is sufficiently slow, the system will remain in the ground state, leading to the ground state of the final Hamiltonian. Therefore, this can be used as a method of computing ground states of spin systems (Ising models). The system is described as
$H_{\text{system}}(t)=A(t)H_{\text{init}}+B(t)H_{\text{Ising}},$ (11)
where $A(t)$ and $B(t)$ define the anneal schedule, $H_{\text{init}}$ is the Transverse field Hamiltonian, and $H_{\text{Ising}}$ is the problem Ising model that we wish to find the optimal solution of. Ising models can easily be converted into QUBOs, and vice versa, and since many optimization problems can be formulated as QUBOs, quantum annealing can be used to sample solutions of a large number of optimization problems [Lucas2014, Yarkoni2021].
@@
$H_{\text{system}}=-\frac{A(s)}{2}\Big{(}\sum_{i}^{p}\sigma^{x}_{i}\Big{)}+\frac{B(s)}{2}\Big{(}H_{\text{Ising}}\Big{)},$ (12)
where $s\in[0,1]$ is a variable called the anneal fraction which defines a weighting of both the transverse field and problem Hamiltonians (by $A(s)$ and $B(s)$) at each point in time during the anneal, $\sigma^{x}_{i}$ is the Pauli matrix acting on each qubit $i$, and $H_{\text{Ising}}$ is the problem Ising model that we wish to sample low energy states from. In our case, $H_{\text{Ising}}$ is the QUBO of eq. (5) that has been converted to an Ising model such that all of the variable states are spins. The user can program a variety of additional features on the device, including different anneal schedules and annealing times.
@@ 2.4.1 Parallel Quantum Annealing Minor Embeddings
In order to make more effective use of the large hardware graph of D-Wave’s Advantage_system4.1 (compared to the QUBO models which are being embedded), we employ parallel quantum annealing, also tiling, in which multiple Ising models can be embedded on the chip such that they are disjoint and are solved in parallel during a single anneal [Pelofske2022parallel, Pelofske2022solving, PhysRevA.91.042314]. As shown in Figure 7, we use multiple all-to-all minor embeddings onto the D-Wave hardware graph, thus allowing us to solve the same problem multiple times simultaneously during a single anneal readout cycle. This allows us to obtain more samples per anneal, but also mitigates local biases and errors that may be present on some parts of the chip (as well as differences in the random minor-embeddings). Moreover, parallel quantum annealing partially mitigates the effects of using a single sub-optimal embedding due to multiple different random embeddings being used, and also different parts of the hardware graph being used (since the D-Wave hardware has non-uniform noise properties across the different components). Note that parallel quantum annealing is also referred to as tiling in the D-Wave documentation  [^2] https://dwave-systemdocs.readthedocs.io/en/samplers/reference/composites/tiling.html. Throughout all experiments, we consider all-to-all embeddings for sizes $64$ and $128$ variables. Those (random) minor embeddings can be computed with the heuristic tool minorminer [minorminer], applied iteratively to the hardware graph in order to obtain disjoint minor embeddings, using the methods described in [Pelofske2022parallel] which attempts to maximize the number of disjoint minor embeddings that can be embedded on the D-Wave QPU hardware graph.
@@
2.4.2 Choice of D-Wave Hardware Parameters
@@
For all experiments involving reverse annealing, the schedule used is $\{[0,1],[10,s],[90,s],[100,1]\}$, where each tuple defines a point in time (from the start to the end of the anneal process) and an anneal fraction $s$. The anneal fraction is the normalized time used to control how the quantum Hamiltonian is moved from the initial superposition of states to the problem QUBO (which is a classical Hamiltonian) during the anneal [annealing]. The anneal schedule is constructed by linear interpolation between those four points. The reverse anneal schedule we use is symmetric with a pause duration of $80$ microseconds. It has an increasing and decreasing ramp on either side of a pause of duration $10$ microseconds. We vary the anneal fraction $s$ at which the pause occurs.
@@ 2.4.3 Quantum Evolution Monte Carlo with Reverse Annealing
We will use the D-Wave quantum annealer to compute a Monte Carlo chain of reverse quantum anneal sequences, in which the best solution of an initial set of anneal-readout cycles is encoded as the initial state of the next anneal, and then the best solutions found from each iteration are used as the initialized classical spin state of the subsequent (reverse) anneals. This process works on the logical problem, meaning after unembedding of all chained qubits.
@@
To be precise, a sequence of reverse anneals are chained together in a Monte Carlo like process, where each subsequent round of reverse anneals is initialized with a classical state that is defined by the best solution found at the last set of reverse anneals. This chain of reverse anneals is initialized with the best solution found from a $100$ microsecond forward anneal with $1000$ samples (qubit measurements) and the standard linear anneal schedule (forward anneal denotes that it is a standard annealing process, initialized in a uniform superposition and not in a specific configuration of spins). Because of the use of parallel quantum annealing, there is a multiplier on the total number of samples obtained at each QEMC step; for the $64$ variable QUBO problems we actually obtain $7,0000$ samples per $1,000$ sample anneal readout cycle, and for the $128$ variable QUBO problems we obtain $2,000$ samples per $1,000$ sample anneal readout cycle. Each reverse annealing step in the chain uses the parameter reinitialize_state=True when executing on the D-Wave quantum annealers, which re-initializes the state of the reverse anneal after each anneal-readout cycle. This technique has been used in other contexts, and is referred to as both “iterative reverse annealing” [Yamashiro2019, Bando2022, Arai2021, pelofske2023initial] and “quantum evolution Monte Carlo” (QEMC) [King2021magnetization, King2018, King2021scaling, Kairys2020, LopezBezanilla2023]. The technique we use here differs slightly from some of these approaches in that we do not use a single anneal-readout cycle at each step of the Monte Carlo chain process, instead we use a batch of anneals at each step and initialize the next anneal with the lowest energy solution found in the current batch of anneals.
Figure 2: Reference fMNIST [xiao2017fashionmnist] image that will be reconstructed using binary sparse coding.
@@
Figure 3: Binary sparse coding image reconstructions of the original image shown in Figure 2. Top left: The best (lowest energy) reconstruction found using D-Wave with standard forward annealing (with no parallel QA and an annealing time of $20$ microseconds). Bottom left: reverse quantum annealing chain of Monte Carlo-like iterations (QEMC). Top right: Single execution of neuromorphic computing on Loihi 2. Bottom right: Iterated warm starting neuromorphic computing with Loihi 2. Combined, the bottom row of plots show that the two iterative approaches yield better solution quality than that non iterative improvement methods shown in the top row. All experimentally computed figure reconstructions are the best mean energy and best mean sparsity across the $16$ QUBO models (for the best parameter combination found for each device or technique).
@@
Figure 4: Iterated warm starting on results on Loihi 2 for $100$ iterations for each of the $16$ QUBO problems (each QUBO contains $64$ variables). Each panel shows the results for one of the $16$ QUBO problems. The minimum energy sample from the converged simulation is plotted at each iteration step.
@@ 2.6 Exact QUBO Solutions Using CPLEX
In order to determine how well all of the evaluated heuristic methods worked at sampling the optimal solution(s) of the generated QUBO models, we used the classical CPLEX [CPLEX] solver which can solve the generated QUBOs exactly. CPLEX is a commercial mathematical optimization software. To this end, we first transform each QUBO into a Mixed Integer Quadratic Programming (MIQP) problem, and solve those problems exactly using the CPLEX solver [CPLEX]. CPLEX deterministically finds and then verifies the optimal solution of a combinatorial optimization problem. Software such as CPLEX are used in the context of either generating good heuristic solutions within a compute time budget, or using as much compute time as is required to verify optimality of a solution – in our case we use the later (i.e., no time limit) in order to compute the optimal energy (ground state energy) for these QUBO problems. CPLEX, although not as performant as new optimization software such as Gurobi, is a very standard classical benchmark for (general) classical optimization techniques. The CPLEX solver does not provide information on whether the QUBOs have multiple optimal solutions, or only $1$, however it is expected that in general these QUBOs will not have degenerate ground states. We leave further investigations on the ground state degeneracy of these binary sparse coding QUBO problems to future research.
@@ 3 Results
Figure 4 presents the iterated warm started learning on Loihi 2 for all $16$ of the $64$ decision variable QUBO models. Interestingly, this process of warm starting the simulation using the state found during the previous computation seems to converge fairly fast to a local minimum for most of the QUBO problems. In these plots, each iteration is showing the minimum energy obtained from a set of $300$ samples obtained during the $6000$ time step simulation (at intervals of $20$ time steps). Notably, $4$ of the plots show high variability of the learning curve for those QUBOs, whereas for most of the other QUBOs there is very little variability of the learning curve. Figure 4 also plots the minimum energies computed using CPLEX. These results are the first that we are aware of for using iterated warm starting on a spiking neuromorphic processor.
@@
Figure 5 shows the QEMC, or iterated reverse annealing, progression when sampling all of the $16$ sparse coding QUBOs using $100$ iterations and varying anneal fractions at which the pause occurred. At each QEMC step, the minimum energy (i.e., the objective function evaluation) sample is plotted. Figure 5 shows that there are some anneal fractions ($s$) that produced the desired monotonic decrease of minimum computed energy at each QEMC step, whereas other anneal fractions did not consistently improve on solution quality. Specifically, the anneal fraction of $s=0.44$ clearly introduced too much of the transverse field Hamiltonian into each simulation, resulting in a loss of information on the best previous obtained solution. On the other hand, the choice $s=0.6$ did not utilize enough of the quantum fluctuations provided by the transverse field Hamiltonian to search for lower energy states, although the improvement was always monotonic. The anneal fractions of $s=0.5$ and $0.55$ showed the best tradeoff in our experiments, allowing for reasonably good exploration of the problem search space. Figure 5 shows that the RA QEMC method was able to find the optimal solution for some, but not all of of the QUBO problems. This is examined in more detail in Table 1. Figure 5 plots the minimum energy computed by simulated annealing as black horizontal dashed lines, and Table 1 shows that those solutions are also optimal as verified by CPLEX. When executing the QEMC simulations of Figure 5, the initial state was chosen to be the best sample found from an initial standard linear schedule forward anneal; that initialization is then repeated with independent anneals for each run of QEMC, meaning that the initial states all have slightly different objective values. If there were multiple solutions with the same objective function cost, one was chosen arbitrarily as the initial vector.
@@
The ordering of the sub-plots (in terms of the QUBO problem being solved) in Figure 4 and Figure 5 is consistent between the two figures, thus facilitating a comparison. Notably, all of the iterated reverse annealing plots of Figure 5 show very clear and consistent convergence behavior (e.g., for $s=0.5$), but the iterated warm starting on Loihi 2 does not have as clear convergence behavior as a function of the iteration index.
@@
Table 1 also reports the sparsity of the optimal bitstring computed by CPLEX. Simulated annealing always correctly found the optimal solution. Loihi 1 found the optimal solution for 00 out of the $16$ QUBOs, so the optimal solution count is not shown. The Loihi 1 results are shown for comparison against Loihi 2, and are taken from the simulations in [henke2023sampling], where the minimum energies are taken from the distribution of $2000$ samples computed across a varied set of device parameters. The minimum energies reported both from iterated reverse quantum annealing and iterated warm starting on Loihi 2 are the minimum energies found from the entirety of the simulation (which, for converged simulations is the last point found after the $100$ iterations). The reported minimum energies from Loihi 2 are taken from the distribution of $100$ samples obtained from the iterated warm starting procedure.
@@
Table 1 verifies that the trained sparse coding QUBOs indeed have solution vectors that are reasonably sparse. However, although unlikely to occur for our types of QUBO models, there could exist multiple variable assignments that give the same optimal objective function value (i.e., degenerate ground states). Moreover, as CPLEX only returns one optimal solution, the reported CPLEX results do not capture the case of more than one optimal solution with different sparsity. We do not report compute time in Table 1 because it is relatively stable across different problem instances: the mean QPU time to obtain $1000$ samples on the D-Wave quantum processor is $1.3$ seconds, the mean wall-clock time to obtain $300$ samples on the Loihi 2 processor (this is $300$ samples taken during a single simulation sweep where we are measuring every $20$ time steps within the total simulation duration of $6000$ time steps) is $6.7$ seconds for the randomly initialized first iteration, and is $5.4$ seconds for each subsequent warm started iteration.
@@
Figure 5: Quantum Evolution Monte Carlo with reverse annealing convergence. Each sub-plot shows the QEMC convergence for each of the $16$ separate, $64$-variable, QUBOs that are the set of patch sub-problems. The anneal fraction (and thereby the transverse field proportion relative to the diagonal problem Hamiltonian) at which the reverse anneal chain is paused is varied, and the resulting minimum energy found in the $1000$ samples at each Monte Carlo step is plotted. All quantum annealing computations used an annealing time of $100$ microseconds.